# Information theory for data-driven model reduction in physics and biology

### v3

@@ IX Methods
Encoder in terms of transfer operator eigenfunctions
To connect the optimal encoder to the transfer operator, we first rewrite (13) in terms of the transition probabilities,
@@
where we have absorbed terms in the exponent which depend only on $h_{t}$ or $x_{t}$ into the normalization factors. Into the above equation, we replace the transition probability with the spectral decomposition
@@
From this, the (7) of the main text immediately follows, where
$f_{n}(h_{t})=\int\text{d}x_{t+\Delta t}\rho_{n}(x_{t+\Delta t})\log p(x_{t+\Delta t}|h_{t})$ (16)
@@
with corrections due to the second eigenfunction given by $f_{2}(h)\approx f_{1}(h)\text{e}^{-\Gamma\Delta t}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t})$ where $\Gamma=\lambda_{1}-\lambda_{2}>0$ denotes the spectral gap. To see this, we compute the Hessian of the IB Lagrangian
$H_{(\mu,n),(\nu,m)}=\frac{\partial^{2}}{\partial f_{n}(h_{\mu})\partial f_{m}(h_{\nu})}\mathcal{L}_{\text{IB}}.$
Here we assume a finite alphabet of size $N_{H}$, i.e. $h_{\mu}$ with $\mu\in\{1,...,N_{H}\}$. At the uniform encoder, i.e. $f_{n}(h)=0$, the Hessian decomposes into a tensor product
$H_{(\mu,n),(\nu,m)}=A^{\beta}_{nm}\otimes G_{\mu\nu}$
where $A^{\beta}$ depends only on the indices of the coefficients and $G$ captures the dependence on $h_{\nu}$. The only part which depends on $\beta$ is $A^{\beta}$.
[…]
Variational IB compared to other dimensionality reduction techniques
@@
Variational IB (VIB) is by no means the only numerical method for performing data-driven model reduction. Here we provide a brief overview of the benefits and shortcomings of VIB with respect to other methods; an extended discussion can be found in the SI.
@@
A second category of non-linear dimensionality reduction techniques are graph-based or similarity-based methods, which typically assume that the data is distributed on a low-dimensional manifold embedded in a higher-dimensional space [95, 96]. One prominent example is diffusion maps [21], which starts from a set of data snapshots and, assuming the system evolves diffusively on short times, constructs an approximate transition matrix from which one can compute eigenfunctions to parameterize the data manifold. The assumption of diffusive dynamics can be violated when data is not sampled sufficiently frequently. This likely explains our finding that VIB produced more well-behaved low-dimensional embeddings on the cyanobacteria dataset (SI Fig. S12). VIB has the additional advantage, relative to this and similar methods, that it explicitly takes dynamics into account without the strong assumptions required by diffusion maps.
@@
In general when investigating a new system it is good practice to start by attempting to perform dimensionality reduction with linear methods such as PCA or DMD because they are fast, straightforward to implement, and easy to interpret. In situations where linear techniques are not sufficient, VIB may be preferable to other methods because it is guaranteed to find dynamically relevant variables (in contrast to diffusion maps, t-SNE, AEs, VAEs, etc.) and it does not require that one performs the carefully tailored preprocessing steps that are required by eDMD or kernel PCA, or other variants of DMD [84, 85, 86, 87]. Additionally, it works well even when the dynamics are highly stochastic as shown in the cyanobacteria dataset.
Acknowledgements
@@ I Background material
The three formulations of the mutual information above are equivalent and useful in diffent settings. For example, formulation as a Kullback-Leibler divergence immediately tells us that mutual information is non-negative.
Information bottleneck
The information bottleneck was originally formulated in Ref. [38] as a rate distortion problem. Rate distortion theory describes how to maximally compress a signal (i.e. minimize its communication rate) such that it remains minimally distorted, which, however, presupposes the knowledge of a distortion function specifying which features need to be preserved[58]. In Ref. [38], Tishby et al. introduce a variant of the problem where the a priori unknown distortion function is not used, but rather one seeks to ensure that the compression retains information about an auxiliary variable $Y$ correlated with the signal. As the correlations with $Y$ define implicitly the relevant features of the signal to be preserved, $Y$ is called the relevance variable. Concretely, we call $X$ the source signal, and $H$ denotes the compressed signal. The random variables form a Markov chain $H\leftrightarrow X\leftrightarrow Y$, meaning that $H$ and $Y$ are conditionally independent given $X$:
$p(y|h,x)$ $=p(y|x)p(h,x)$
$p(h|x,y)$ $=p(h|x)p(y,x)$
$\mathcal{L}_{\text{IB}}[p(h|x)]$ $=I(X,H)-\beta I(Y,H).$ (S2)
$\begin{split}\mathcal{L}_{\text{IB}}[p(h|x)]&=\sum_{x,h}p(h|x)p(x)\log\frac{p(h|x)}{p(h)}-\beta\sum_{y,h}p(y,h)\log\frac{p(y,h)}{p(h)p(y)}+\sum_{x}\lambda(x)\left(1-\sum_{h}p(h|x)\right).\end{split}$ (S3)
$\frac{\delta}{\delta p(h|x)}p(h^{\prime})$ $=\delta(h-h^{\prime})p(x)$
which can be rearranged to give the optimal encoder
$p(h|x)=\frac{p(h)}{N(x)}\exp\left[-\beta D_{KL}(p(y|x)\|p(y|h))\right].$ (S4)
$p(h|x)=\frac{p(h)}{N(x)}\exp\left[\beta\int dy\,p(y|x)\log p(y|h)\right].$ (S5)
When $\beta<1$, it follows from the data processing inequality $I(X,H)\geq I(Y,H)$ that (S2) is minimized by a trivial encoder $p(h|x)=p(h)$. In this case, $\mathcal{L}_{\text{IB}}=0$ and no information passes through the bottleneck. As $\beta$ is increased, more information is allowed through the bottleneck until the relevant variables begin to gain a dependence on the state $x$. This occurs suddenly for a certain value of $\beta=\beta_{1}>1$ at which the encoder becomes non-uniform and $I(H,X)$ becomes non-zero. This is referred to as an IB transition; for increasing $\beta$, there may be a sequence of transitions at $\beta_{2},\beta_{3},$... etc.
@@
The Fokker-Planck equation may be written
Numerically solving “exact” IB and Ulam approximations of the transfer operator
The optimal IB encoder (S4) can be found using the Blahut-Arimoto (BA) algorithm [58]. As described in detail in Refs. [38, 97] and sketched in Fig. S4, the algorithm is an iterative procedure, where the encoder at iteration $k+1$ is updated according to
@@
II Model reduction in general
Consider a dynamical system which can be described by the evolution of the random variable $X_{t}$. In the case where $X_{t}$ is high-dimensional but undergoes “stereotyped” or regular behavior, it often suffices to describe the system with a reduced dynamical variable $H_{t}$, where $H_{t}$ is lower dimensional than $X_{t}$. Schematically, a reduced model would be used in the following way to predict the system’s evolution over a time $\Delta t$, for any choice of starting time $t$:
Figure S1: Two views of model reduction. (a) In one view, the reduced model should act as a useful surrogate of the full system. After reducing the system via an encoder $h_{\text{enc}}$, the reduced state is evolved forward in time via $\mathcal{S}^{\Delta t}$ and subsequently decoded by $h_{\text{dec}}$ to obtain an estimate of the full system $X_{t+\Delta t}$. In general, if significant information has been lost during encoding, the decoding will be probabilistic to produce a distribution of possible future states. (b) Alternatively, one may insist only that the encoding is “consistent” with respect to the dynamics (and is not necessarily informative about the full system). In this picture, it should be equivalent to either encode the future state, or to first encode and then evolve the encoded state.
A reduced description is useful only if it can be used to simulate the system forward in time and make predictions about the future. In other words, one aims to find a reduction $h(X)$ and an evolution rule $\mathcal{S}$ so that the above diagram commutes. Reducing the initial state and evolving it should be equivalent to first evolving the full system, and then reducing the final state.
@@ II Model reduction in general
    $d\big(h_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)$ (S15)
@@
    $d\big(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}),h_{\text{enc}}(X_{t+\Delta t})\big)$ (S16)
@@
In general, one may determine the functions $h_{\textbf{enc}}$, $h_{\textbf{dec}}$ and the operator $\mathcal{S}^{\Delta t}$ as the solution to an optimization problem
$\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}\quad\dim h_{\text{enc}}(X_{t})+\beta_{1}\,d\big(h_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)+\beta_{2}\,d\big(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}),h_{\text{enc}}(X_{t+\Delta t})\big)$ (S17)
The parameters $\beta_{1}$ and $\beta_{2}$ must be chosen to reflect the desired properties of the reduced model. The choice of $\beta_{2}=0$ or $\beta_{1}=0$ correspond to the two views of model reduction described above, but in principle both terms may be present. In both the main text and here, we focus on the cases where only one term is present, but comment briefly on the case $\beta_{1},\beta_{2}\neq 0$ below. It turns out that the case $\beta_{1}=0$ is uninteresting (though see “Relation to other work” below). The reason is that the objective function may then be trivially solved by $h(X_{t})=\text{const}$ and $\mathcal{S}^{\Delta t}=\text{id}$. For any choice of dimension, this will minimize the second term of the objective. Consequently, we take our model reduction objective function to be
$\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}\mathcal{L}_{\text{MR}}\equiv\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}\quad\dim h_{\text{enc}}(X_{t})+\beta_{1}\,d\big(h_{\text{dec}}^{-1}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big).$ (S18)
In the following, we will take the perspective that every step in the diagram above is, or may be, stochastic. That is, the stochastic time evolution of $X_{t}$ and $H_{t}$ are described by stochastic functions $\mathcal{S}$, $\mathcal{T}$, or equivalently by an evolution operator that acts on distributions
$p_{X_{t+\Delta t}}(x)$ $=[U^{\Delta t}p_{X_{t}}](x)$ (S19)
$p_{H_{t+\Delta t}}(h)$ $=[V^{\Delta t}p_{H_{t}}](h).$ (S20)
The encoding map $h(X_{t})$ may also be a stochastic function, whose output is described by a distribution $p(h|x)=p(H_{t}=h|X_{t}=x)$.
@@
$\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}\quad\dim h_{\text{enc}}(X_{t})-\beta_{1}\,I\big(h_{\text{dec}}^{-1}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)$ (S21)
@@
As stated in the main text, this choice of “distance” has dramatic consequences for our optimization problem. To see this, we note that the latent evolution operator $\mathcal{S}$ and decoder $h^{-1}_{\text{dec}}$ only occurs in the second term, so we may write
$\min_{h_{\text{enc}},\mathcal{S},h^{-1}_{\text{dec}}}\mathcal{L}_{\text{MR}}=\min_{h_{\text{enc}}}\left[\dim h_{\text{enc}}(X_{t})-\min_{\mathcal{S},h^{-1}_{\text{dec}}}\beta_{1}\,I\big(h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)\right]$ (S22)
The variables appearing in the above equation the Markov chain
$X_{t+\Delta t}\,\,\text{---}\,\,X_{t}\,\,\text{---}\,\,h_{\text{enc}}(X_{t})\,\,\text{---}\,\,\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})\,\,\text{---}\,\,h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}))$
where $X\,\,\text{---}\,\,Y\,\,\text{---}\,\,Z$ means that $Z$ is conditionally independent of $X$, given $Y$: $p(z,x|y)=p(z|y)p(x|y)$. The data processing inequality states that, given a Markov chain $X\,\,\text{---}\,\,Y\,\,\text{---}\,\,Z$, then
$I(X,Z)\leq I(X,Y).$
In particular, if $Z=g(Y)$ then $I(X,g(Y))\leq I(X,Y)$, where equality is obtained if $g$ is a bijection [58]. Consequently, we have
$I(h^{-1}_{\text{dec}}(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t})$ $\leq I(\mathcal{S}^{\Delta t}h_{\text{enc}}(X_{t}),X_{t+\Delta t})$
$\leq I(h_{\text{enc}}(X_{t}),X_{t+\Delta t}),$
where the inequalities are maximized if $h^{-1}_{\text{dec}}$ and $\mathcal{S}^{\Delta t}$ are both bijective functions. Notably the bound is saturated if $\mathcal{S}^{\Delta t}=\text{id}$ and $h^{-1}_{\text{dec}}=\text{id}$, so that the objective function takes the form
$\min_{h_{\text{enc}}}\mathcal{L}_{\text{MR}}=\min_{h_{\text{enc}}}\left[\dim h_{\text{enc}}(X_{t})-\beta_{1}\,I\left(h_{\text{enc}}(X_{t}),X_{t+\Delta t}\right)\right]$ (S23)
Conceptually, this means that the purely information-theoretic framework we have taken does not tell us how to evolve the reduced variable forward in time, only how to choose the optimal reduced variables. However, the fact that any bijective $\mathcal{S}^{\Delta t}$ is a solution should not be taken to mean that $H_{t}$ has no dynamics, or that its dynamics are undefined. Instead, it simply means that the mutual information is insensitive to the choice of $\mathcal{S}^{\Delta t}$, because this transformation does not impact the information contained about $X_{t+\Delta t}$. The dynamics of $H_{t}$ may be obtained by encoding $X_{t}$ at each time $t$. To determine which $\mathcal{S}^{\Delta t}$ is optimal among all choices of bijective transformations, one must introduce and solve another objective function, for example $\min_{\mathcal{S}}\lVert H_{t+\Delta t}-\mathcal{S}^{\Delta t}H_{t}\rVert^{2}$, which again requires a choice of distance function.
In practice, one may directly solve (S23) by fixing $\dim h_{\text{enc}}$ to various different values, and for each solving for the $h_{\text{enc}}$ which optimizes the mutual information with $X_{t+\Delta t}$. To gain an analytical understanding of what this objective is solving, we instead replace the dimensionality reduction term with a term that similarly restricts the capacity of the reduced variable; namely, the mutual information $I(X_{t},h_{\text{enc}}(X_{t}))$. As a result of this choice, the objective becomes
@@
This is the information bottleneck objective [38, 36] (see Section I). The benefit of working with this objective is that it is analytically tractable, allowing us gain theoretical insight into the form of solutions of (S24) which we expect to be similar to solutions of (S23). Indeed in practice we see that, for fixed $\dim h_{\text{enc}}$, solutions to S24 (which may be found by setting $\beta_{1}\gg 1$) are qualitatively similar to the solution to (S24).
@@
We argued above that the $\beta_{1}=0$ case for the objective function (S17) is uninteresting because of the existence of a trivial solution. However, the situation changes when distances are replaced by mutual information. In that case, the objective becomes
@@
This objective may be thought of as an information bottleneck, but with an additional term $I(H_{t},H_{t+\Delta t})$ that encourages the encoding to learn a variable which is predictive of its own future. We do not consider this case further beyond noting that the addition of such a term does not affect our perturbative analysis below.
III Information theoretic quantities for dynamical systems

### v4

@@ Appendix C A primer on the information bottleneck
$p(h|x)=\frac{p(h)}{N(x)}\exp\left[\beta\int dy\,p(y|x)\log p(y|h)\right].$ (46)
When $\beta<1$, it follows from the data processing inequality $I(X,H)\geq I(Y,H)$ that Eq. (43) is minimized by a trivial encoder $p(h|x)=p(h)$. In this case, $\mathcal{L}_{\text{IB}}=0$ and no information passes through the bottleneck. As $\beta$ is increased, more information is allowed through the bottleneck until the relevant variables begin to gain a dependence on the state $x$. This occurs suddenly for a certain value of $\beta=\beta_{1}>1$ at which the encoder becomes non-uniform and $I(H,X)$ becomes non-zero. This is referred to as an IB transition; for increasing $\beta$, there may be a sequence of transitions at $\beta_{2},\beta_{3},$... etc.
Appendix D The optimal encoder in the limit of high compression
To understand the form of the optimal encoder in terms of transfer operator eigenfunctions, we first rewrite Eq. (46) in terms of the transition probabilities,
@@ Appendix D The optimal encoder in the limit of high compression
where we have absorbed terms in the exponent which depend only on $h_{t}$ or $x_{t}$ into the normalization factors. In the above equation, we replace the transition probability with the spectral decomposition
@@
From this Eq. (16) of the main text immediately follows, where
$f_{n}(h_{t})=\int\text{d}x_{t+\Delta t}\rho_{n}(x_{t+\Delta t})\log p(x_{t+\Delta t}|h_{t}),$ (49)
@@
with corrections due to the second eigenfunction given by $f_{2}(h)\approx f_{1}(h)\text{e}^{-\Gamma\Delta t}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t})$ where $\Gamma=\lambda_{1}-\lambda_{2}>0$ denotes the spectral gap. Here we assume a finite alphabet of size $N_{H}$, i.e. $h_{\mu}$ with $\mu\in\{1,...,N_{H}\}$, but in the SI we consider the case where $p(h|x)$ is Gaussian which is relevant for variational IB. To prove the above, we compute the Hessian of the IB Lagrangian
$H_{(\mu,n),(\nu,m)}$ $=\frac{\partial^{2}}{\partial f_{n}(h_{\mu})\partial f_{m}(h_{\nu})}\mathcal{L}_{\text{IB}}$
$=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\beta\left\langle\langle\tilde{\phi}_{n}\rangle_{p(\cdot|y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot|y)}\right\rangle_{y}-(1-\beta)\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$
where $\tilde{\phi}_{n}(x)=\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x)$. We see that $H$ has the form of a Kronecker (tensor) product
$H_{(\mu,n),(\nu,m)}$ $=A^{\beta}_{nm}\otimes G_{\mu\nu}.$
Eigenvalues of $H$ are given by products of eigenvalues of $A^{\beta}$ and $G$, while eigenvectors are given by the tensor product of eigenvectors of $A^{\beta}$ and $G$. Note that $G$, which is nothing other than the covariance matrix for a multinomial distribution, does not depend on $\beta$. Therefore we expect the behavior at the transition to be governed by $A^{\beta}$.
We are concerned with the sign of the eigenvalues of $H$. A negative eigenvalue indicates that $\mathcal{L}_{\text{IB}}$ is unstable to a perturbation in $f$, which means the loss can be lowered by changing $f$ away from the trivial encoder at $f_{n}=0$. Because the eigenvalues of a tensor product of matrices are products of the eigenvalues of the component matrices, the eigenvalues of $H$ change sign when those of $A^{\beta}$ change.
For equilibrium systems, i.e. those with no steady state flux, the matrix is given by
$A_{nm}=\delta_{nm}\text{e}^{2\lambda_{n}\Delta t}\left(1-\beta\text{e}^{2\lambda_{n}\Delta t}\right)$ (51)
For non-equilibrium systems, we consider the $\times2$ block of $A$ to study the effect of corrections due to the second transfer operator eigenvector:
$A$ $=\begin{pmatrix}A_{2\times 2}&0\\ 0&0\end{pmatrix}+\mathcal{O}(\text{e}^{\lambda_{3}\Delta t})$
The stability of $A$ is given to the desired order by the stability of $A_{2\times 2}$. As we show in the SI Section S4, the first eigenvector of $A$ to change sign takes the form $(f_{1},0)$ with a correction of size $\mathcal{O}\left(\langle\phi_{1}\phi_{2}\rangle\text{e}^{-\Gamma\Delta t}\right)$, which is controlled both by the size of the spectral gap ($\Gamma>0$) and the amount of overlap between $\phi_{1}$ and $\phi_{2}$.
The critical values of $\beta$ do not depend on the dimensionality of the latent variable $N_{H}$ [127, 128]. Therefore, because we are interested in only the first perturbation from the uniform encoder, we may consider the case $N_{H}=2$. As we show in the SI for both this discrete case and the case of a Gaussian encoder, the transition is continuous, so that the encoder depends only on $f_{1}$ after the transition, albeit weakly.
Appendix E Variational IB compared to other dimensionality reduction techniques
@@
Here we provide a brief overview of the benefits and shortcomings of variational IB (VIB) with respect to other methods, illustrated on the cyanobacteria dataset in Fig. 8. An extended discussion can be found in the SI.
Figure 8: Comparison of various dimensionality reduction techniques on cyanobacteria dynamics. We compare various dimensionality reduction techniques in the case study of cyanobacteria dynamics. As illustrated on the top row, in this case a good dimensional reduction should produce trajectories in latent space that separate well the two regimes of low and high synchronization, that encode the (noisy) cyclic nature of the dynamics, and if possible that encode the amount of synchronization in the size (radius) of the cycle. For each method, we plot the trajectories of all cyanobacteria colonies when projected onto the dominant modes (whose definition depends on the particular method). Blue trajectories show the highly synchronized colonies corresponding to high theophylline concentration, black/gray trajectories show the unsynchronized colonies from low theophylline conditions. Unlike the first four methods, deep neural networks (bottom right) directly produce a two-dimensional latent variable and do not provide access to subleading modes. (a) Principal component analysis reduces the dimensionality of data by finding a projection onto the directions of most variance (sketch). (Top) Projection of the cyanobacteria trajectories onto several pairs of principal components, $a_{i}(t)=\mathbf{x}(t)\cdot\mathbf{m}_{i}$. (Bottom) The first four principal components. Here, the state is composed of two time-lagged images of the cyanobacteria colony; the top half of $\mathbf{m}$ corresponds to the earlier image, the bottom half corresponds to the lagged image. (b) Dynamic mode decomposition is performed on pairs of data points and aims to find a linear evolution operator $\mathbf{A}$ that links them (sketch). (Right) Trajectories of the projection of the state onto several eigenvectors of $\mathbf{A}$. The eigenvectors themselves are shown below (the imaginary parts of the first two modes are uniformly zero). (c) Independent component analysis seeks a linear projection of the data onto statistically independent components $\mathbf{s}_{i}$. In other words, the distribution $p(s_{i}|s_{j})$ is independent of $s_{j}$ (sketch). (Right, top) Trajectories of the projection of the state onto several independent components. (Bottom) columns of the mixing matrix; the components $s_{i}$ describe how the weighting of these columns evolves in time. (d) (Left) Diffusion maps build a graph of the observed state variables with edges weights determined by their distance from one another. (Right) Projection of the trajectories onto the first non-trivial eigenfunctions (because $a_{0}=$const) of the inferred transfer operator on the graph. (e) A (time-lagged) autoencoder is a deep neural network architecture which encodes the state $\mathbf{x}_{t}$ into a low-dimensional latent space. From the encoding one then tries to reconstruct the future state. (Right) Trajectories of the cyanobacterial colonies when encoded into a two-dimensional latent space. (f) A variational autoencoder is similar to a standard autoencoder, where one instead learns a distribution over possible encodings, $p(\mathbf{h}_{t}|\mathbf{x}_{t})$. (Right) Trajectories of the cyanobacterial colonies when encoded into a two-dimensional latent space. (g) Implementation of the network in Ref. [49] which encodes the state using a time-lagged autoencoder with an additional loss term that aims to enforce linearity of the embedded dynamics, $\mathbf{h}_{t+\Delta t}\approx\mathbf{K}\mathbf{h}_{t}$. (h) Latent trajectories produced by variational IB (VIB).
E.1 Linear methods
@@
E.2 Nonlinear methods
A second category of non-linear dimensionality reduction techniques are graph-based or similarity-based methods, which typically assume that the data is distributed on a low-dimensional manifold embedded in a higher-dimensional space [134, 135]. One prominent example is diffusion maps [15], which starts from a set of data snapshots and, assuming the system evolves diffusively on short times, constructs an approximate transition matrix from which one can compute eigenfunctions to parameterize the data manifold. The assumption of diffusive dynamics can be violated when data is not sampled sufficiently frequently. This likely explains our finding that VIB produced more well-behaved low-dimensional embeddings on the cyanobacteria dataset (Fig. 8). VIB has the additional advantage, relative to this and similar methods, that it explicitly takes dynamics into account without the strong assumptions required by diffusion maps.
E.3 Deep neural network based methods
@@
E.4 Summary
When investigating a new system, we suggest to start by attempting to perform dimensionality reduction with linear methods such as PCA or DMD because they are fast, straightforward to implement, and easy to interpret. In situations where linear techniques are not sufficient, VIB may be preferable to other methods because it is guaranteed to find dynamically relevant variables (in contrast to diffusion maps, t-SNE, AEs, VAEs, etc.) and it does not require that one performs the carefully tailored preprocessing steps that are required by eDMD or kernel PCA, or other variants of DMD [110, 111, 112, 113]. Additionally, it works well even when the dynamics are highly stochastic, as we have illustrated with the cyanobacteria dataset.
@@ E.4 Summary
The three formulations of the mutual information above are equivalent and useful in different settings. For example, formulation as a Kullback-Leibler divergence immediately tells us that mutual information is non-negative.
@@
Solving IB exactly and Ulam approximations of the transfer operator
The optimal IB encoder Eq. (45) can be found using the Blahut-Arimoto (BA) algorithm [82]. As described in detail in Refs. [41, 52] and sketched in Fig. S1, the algorithm is an iterative procedure, where the encoder at iteration $k+1$ is updated according to
@@
S2 Dimensionality reduction through the lens of information theory
Consider a dynamical system which can be described by the evolution of the random variable $X_{t}$. In the case where $X_{t}$ is high-dimensional but undergoes stereotyped or regular behavior, it often suffices to describe the system with a reduced dynamical variable $H_{t}$, where $H_{t}$ is lower dimensional than $X_{t}$. Schematically, a reduced model would be used in the following way to predict the system’s evolution over a time $\Delta t$, for any choice of starting time $t$:
Figure S2: Two views of model reduction. (a) In one view, the reduced model should act as a useful surrogate of the full system. After reducing the system via an encoder $h_{\text{enc}}$, the reduced state is evolved forward in time via $\mathcal{V}^{\Delta t}$ and subsequently decoded by $h_{\text{dec}}^{-1}$ to obtain an estimate of the full system $X_{t+\Delta t}$. In general, if significant information has been lost during encoding, the decoding will be probabilistic to produce a distribution of possible future states. (b) Alternatively, one may insist only that the encoding is “consistent” with respect to the dynamics (and is not necessarily informative about the full system). In this picture, it should be equivalent to either encode the future state, or to first encode and then evolve the encoded state.
A reduced description is useful only if it can be used to simulate the system forward in time and make predictions about the future. In other words, one aims to find a reduction $h(X)$ and an evolution rule $\mathcal{V}$ so that the above diagram commutes. Reducing the initial state and evolving it should be equivalent to first evolving the full system, and then reducing the final state.
@@
    $d\big(h_{\text{dec}}^{-1}(\mathcal{V}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)$ (S11)
@@
    $d\big(\mathcal{V}^{\Delta t}h_{\text{enc}}(X_{t}),h_{\text{enc}}(X_{t+\Delta t})\big)$ (S12)
@@
In general, one may determine the functions $h_{\text{enc}}$, $h_{\text{dec}}$ and the operator $\mathcal{V}^{\Delta t}$ as the solution to an optimization problem
$\min_{h_{\text{enc}},\mathcal{V},h^{-1}_{\text{dec}}}\quad\dim h_{\text{enc}}(X_{t})+\beta_{1}\,d\big(h_{\text{dec}}^{-1}(\mathcal{V}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)+\beta_{2}\,d\big(\mathcal{V}^{\Delta t}h_{\text{enc}}(X_{t}),h_{\text{enc}}(X_{t+\Delta t})\big)$ (S13)
The parameters $\beta_{1}$ and $\beta_{2}$ must be chosen to reflect the desired properties of the reduced model. The choice of $\beta_{2}=0$ or $\beta_{1}=0$ correspond to the two views of model reduction described above, but in principle both terms may be present. In both the main text and here, we focus on the cases where only one term is present, but comment briefly on the case $\beta_{1},\beta_{2}\neq 0$ below. It turns out that the case $\beta_{1}=0$ is uninteresting. The reason is that the objective function may then be trivially solved by $h(X_{t})=\text{const}$ and $\mathcal{V}^{\Delta t}=\text{id}$. For any choice of dimension, this will minimize the second term of the objective. Consequently, we take our model reduction objective function to be
$\min_{h_{\text{enc}},\mathcal{V},h^{-1}_{\text{dec}}}\mathcal{L}\equiv\min_{h_{\text{enc}},\mathcal{V},h^{-1}_{\text{dec}}}\quad\dim h_{\text{enc}}(X_{t})+\beta_{1}\,d\big(h_{\text{dec}}^{-1}(\mathcal{V}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big).$ (S14)
@@
$\min_{h_{\text{enc}},\mathcal{V},h^{-1}_{\text{dec}}}\quad\dim h_{\text{enc}}(X_{t})-\beta_{1}\,I\big(h_{\text{dec}}^{-1}(\mathcal{V}^{\Delta t}h_{\text{enc}}(X_{t})),X_{t+\Delta t}\big)$ (S15)
@@
As stated in the main text, this choice of “distance” has dramatic consequences for our optimization problem. As we show in Appendix C in the main text, the objective then simplifies to
$\min_{h_{\text{enc}}}\mathcal{L}=\min_{h_{\text{enc}}}\left[\dim h_{\text{enc}}(X_{t})-\beta_{1}\,I\left(h_{\text{enc}}(X_{t}),X_{t+\Delta t}\right)\right]$ (S16)
Conceptually, this means that the purely information-theoretic framework we have taken does not tell us how to evolve the reduced variable forward in time, only how to choose the optimal reduced variables. However, the fact that any bijective $\mathcal{V}^{\Delta t}$ is a solution should not be taken to mean that $H_{t}$ has no dynamics, or that its dynamics are undefined. Instead, it simply means that the mutual information is insensitive to the choice of $\mathcal{V}^{\Delta t}$, because this transformation does not impact the information contained about $X_{t+\Delta t}$. The dynamics of $H_{t}$ may be obtained by encoding $X_{t}$ at each time $t$. To determine which $\mathcal{V}^{\Delta t}$ is optimal among all choices of bijective transformations, one must introduce and solve another objective function, for example $\min_{\mathcal{V}}\lVert H_{t+\Delta t}-\mathcal{V}^{\Delta t}H_{t}\rVert^{2}$, which again requires a choice of distance function.
In practice, one may directly solve Eq. (S16) by fixing $\dim h_{\text{enc}}$ to various different values, and for each solving for the $h_{\text{enc}}$ which optimizes the mutual information with $X_{t+\Delta t}$. To gain an analytical understanding of what this objective is solving, we instead replace the dimensionality reduction term with a term that similarly restricts the capacity of the reduced variable; namely, the mutual information $I(X_{t},h_{\text{enc}}(X_{t}))$. As a result of this choice, the objective becomes
@@
This is the information bottleneck objective [41, 39] (Appendix C in the main text). The benefit of working with this objective is that it is analytically tractable, allowing us gain theoretical insight into the form of solutions of Eq. (S17) which we expect to be similar to solutions of Eq. (S16). Indeed in practice we see that, for fixed $\dim h_{\text{enc}}$, solutions to S17 (which may be found by setting $\beta_{1}\gg 1$) are qualitatively similar to the solution to Eq. (S17).
@@
We argued above that the $\beta_{1}=0$ case for the objective function Eq. (S13) is uninteresting because of the existence of a trivial solution. However, the situation changes when distances are replaced by mutual information. In that case, the objective becomes
@@
This objective may be thought of as an information bottleneck, but with an additional term $I(H_{t},H_{t+\Delta t})$ that encourages the encoding to learn a variable which is predictive of its own future. We do not consider this case further beyond noting that the addition of such a term does not affect our perturbative analysis in Section S4.
S3 Information-theoretic quantities for dynamical systems